# LOGCAN++: LOCAL-GLOBAL CLASS-AWARE NETWORK FOR SEMANTIC SEGMENTATION OF REMOTE SENSING IMAGES

### v1

@@ Abstract
Remote sensing images usually characterized by complex backgrounds, scale and orientation variations, and large intra-class variance. General semantic segmentation methods usually fail to fully investigate the above issues, and thus their performances on remote sensing image segmentation are limited. In this paper, we propose our LOGCAN++, a semantic segmentation model customized for remote sensing images, which is made up of a Global Class Awareness (GCA) module and several Local Class Awareness (LCA) modules. The GCA module captures global representations for class-level context modeling to reduce the interference of background noise. The LCA module generates local class representations as intermediate perceptual elements to indirectly associate pixels with the global class representations, targeting at dealing with the large intra-class variance problem. In particular, we introduce affine transformations in the LCA module for adaptive extraction of local class representations to effectively tolerate scale and orientation variations in remotely sensed images. Extensive experiments on three benchmark datasets show that our LOGCAN++ outperforms current mainstream general and remote sensing semantic segmentation methods and achieves a better trade-off between speed and accuracy.
@@ I Introduction
Semantic segmentation of remote sensing images, aims to assign explicit categories to each pixel of a geospatial object in an image. This technique has a wide range of applications in fields such as environmental protection [1], resource investigation [2], and urban planning [3, 4]. In recent years, the rapid development of remote sensing platforms and sensors has resulted in higher spatial, spectral and temporal resolution of remote sensing images [5]. This provides the possibility of detailed observation of the Earth’s surface features. As a result, more advanced semantic segmentation techniques for remote sensing images are constantly developed to meet the needs to process rich data.
@@
In this paper, we focus on designing a effective network structure to address the aforementioned issues. We first analyze the convolution-based neural networks. Existing convolutional neural network (CNN)-based semantic segmentation methods focus on context modeling [6, 7, 8, 9, 10], which can be divided into spatial context modeling and relational context modeling. Spatial context modeling methods, such as PSPNet [6] and DeepLabv3+ [11], use spatial pyramid pooling (SPP) or empty spatial pyramid pooling (ASPP) to integrate spatial context information. Although these methods can capture homogeneous contextual dependencies, they ignore the differences between classes, which may produce unreliable contextual information when dealing with remote sensing images with complex features and large spectral differences. Relational context modeling methods capture heterogeneous dependencies based on attention mechanism. nonlocal neural network [7] is proposed to compute the similarity of pixel pairs within an image for weighted aggregation, while DANet [8] uses spatial attention and channel attention for selective aggregation. However, such intensive attention operations are not applicable to remote sensing images, as their complex backgrounds would cause these operations to generate a large amount of background noises, leading to degraded semantic segmentation performances. Although recently proposed class-level context modeling approaches (e.g., ACFNet [12], OCRNet [13] and HMANet [14]) can reduce background noise interference caused by dense attention by integrating class-level contexts (i.e., class centers) using global class representations, they are still not perform well in remote sensing images. This is caused by the fact that remote sensing images have large intra-class variance, which leads to a large semantic gap between features and global class centers, and thus hindering models’ performances.
@@
  • We introduce the affine transformation for the first time to extract local class centers to adapt to geospatial objects with different sizes, shapes and orientations. This can cope well with the scale and orientation variations of remote sensing images.
@@
This paper is an extension of our previous conference paper [20]. This work extends the previous work in three aspects. 1) We introduce affine changes to improve the LCA module. It copes well with scale and orientation variations of remotely sensed images by converting the default local window to a target quadrilateral to accommodate geospatial objects of different sizes, shapes, and orientations. 2) We extend the attention operation in the LCA module to be multi-head. so as to improve its ability to capture complex pixel-class-center relationships and to help avoid the overfitting problem that may result from a single attention mechanism. 3) We standardized the experimental setup and added sufficient experiments. We first added experiments on the larger LoveDA dataset. In addition, we provide more quantitative comparison results, qualitative analysis, and ablation experiments. We demonstrate the effectiveness of LOGCAN++ with more adequate experimental data.
@@ II-A General semantic segmentation
In summary, the above methods have achieved remarkable results on natural images. However, when applied to remotely sensed images, they perform unsatisfactorily by ignoring the characteristics of remote sensing images with complex backgrounds, scale and orientation variations, and high intra-class variance.
@@ II-B Semantic segmentation of remote sensing images
Recently, segmentation models for generic application scenarios (e.g., LANet[55], MANet [56], PointFlow[57], SCO[58], GLOTs [59] and Farseg++[60]) also have been widely investigated. However, these methods still fail to fully consider the characteristics of remote sensing images with complex backgrounds, scale and orientation variations, and high intra-class variance, limiting the further improvement of model performance. For example, Ding et al. [55] proposes to enhance the embedding of contextual information using local attention blocks, which works to mitigate object scale variation and intra-class variance but ignores the challenges posed by complex backgrounds. Zheng et al. [61, 60] enhances the differentiation of foreground features by learning foreground-scene relations to associate foreground-related contexts. Although it enhances the features in the foreground well, it ignores the large intraclass variance in remotely sensed images, leading to poorer intraclass feature compactness. Especially, these methods have insufficient consideration for the characteristics of scale and orientation variations in remote sensing images. Therefore, the performance of these methods still needs to be improved due to the deficiencies in structural design.
@@ III-C Local Class Aware Module
After that, we implement the interaction between features and class centers using the multi-head cross-attention operation. Specifically, for each head $H_{i}$, the corresponding affinity matrix $R_{r}^{i}$, which represents the similarity between the pixel and the local class representations, is obtained as follows,
@@ III-D Affine Transformation Block
The ideal local class center should be the average semantic feature of objects in the local area. However, due to the Satellite sensor’s unique overhead view and large imaging range, objects in remotely sensed images usually present different orientations and possess significant scale differences. Fixed local windows often do not model these objects well. Therefore, we propose an affine transformation block (ATB) to extract higher quality local class centers, where the fixed local window is converted to a target quadrilateral to accommodate geospatial objects with different sizes, shapes, and orientations.
@@
Then, we obtain the scaling factor $\Psi$, rotation factor $\Theta$ and offset factor $\Delta$ by feeding the obtained local contextual feature $G_{l}$ linear mapping and LeakyReLU activation function as:
$\Theta,\Delta,\Psi=\text{LeakyReLU}(\phi(G_{l})).$ (7)
Here, we give a brief description of the affine transformation function. As shown in the lower right of Fig. 3, given an initial local window, $x_{l}$, $y_{l}$ and $x_{r}$, $y_{r}$ denote the coordinates of the upper-left and lower-right corners of the initial window, respectively, while $x_{c}$, $y_{c}$ denote the coordinates of the window center. Then, its window transformation process is defined as:
$\begin{bmatrix}\hat{x}_{l/r}\\ \hat{y}_{l/r}\end{bmatrix}=\begin{bmatrix}x_{c}\\ y_{c}\end{bmatrix}+\begin{bmatrix}\Delta_{x}\\ \Delta_{y}\end{bmatrix}+\begin{bmatrix}\cos\Theta&\sin\Theta\\ -\sin\Theta&\cos\Theta\end{bmatrix}\begin{bmatrix}|x_{l/r}-x_{c}|\cdot\Psi\\ |y_{l/r}-y_{c}|\cdot\Psi\end{bmatrix}.$ (8)
We follow this formula to scale, rotate and offset the initial window to obtaind transformed local featurte $\hat{R}_{l}$ and corresponding pre-classified distribution $\hat{D}_{l}$. Since these factors are learned based on the pixel features, the transformed local window can be adaptively adjusted according to the size and orientation of the feature object. As a result, our ATB can help to cope with the scale and orientation changes in the remote sensing image well.
@@ IV-B Evaluation Metrics
IV-C Implementation Details
V EXPERIMENTAL RESULTS AND ANALYSIS
V-A Results on the ISPRS Vaihingen dataset
V-A1 Quantitative analysis
To validate the effectiveness of our LOGCAN++, we first carry out experiments on the Vaihingen dataset. The comparison methods include convolution-based methods such as PSPNet [6], DeepLabV3+ [11], Semantic FPN [63], DANet [8], OCRNet [13], FarSeg [61], transformer-based methods such as segmenter [15], DC-Swin [64], UNetFormer [17], EfficientViT [65] and diffusion based methods such as DDP [66]. As shown in Table I, our LOGCAN++ achieved the optimal segmentation performance. Specifically, compared to recent general segmentation methods such as DDP and RS segmentation methods such as UNetFormer, we have an increase of 2.05% and 1.23% in mIoU, respectively. In particular, for classes with large variance such as buildings, our approach achieved a mIoU value of 91.85%, which outperforms other mainstream methods, validating the effectiveness of our local global class aware strategy. In addition, for small objects such as car, we have a 1.82% mIoU improvement compared to DDP. This indicates that the introduction of affine transformations enhanced the fine-grained recognition of small objects. In addition, although we have a slightly higher number of parameters and FLOPs than Semantic FPN, we still have an advantage in model efficiency compared to other methods. Overall, the proposed LOGCAN++ achieve a better trade-off between accuracy and efficiency.
TABLE II: Comparison with state-of-the-art methods on the test set of the ISPRS Postdam dataset. Per-class best performance is marked in bold. The FLOPs is calculated with the size of $512\times 512\times 3$. The best value in each column is bolded.
@@ V-A2 Qualitative analysis
In addition, Fig. 5 visualizes the segmentation masks output by different models to qualitatively compare the segmentation performance of our LOGCAN++ with these competitors, where all input images come from the Vaihingen test set. For the first and second images, buildings have more complete shapes and clearer boundaries in the output mask of LOGCAN++. For the third image, our LOGCAN++ is able to recognize cars accurately, while other methods such as UNetFormer and DDP misidentify them as buildings. These results show that our LOGCAN++ has better semantic recognition ability and visualization performance.
V-B Results on the ISPRS Potsdam dataset
V-B1 Quantitative analysis
V-B2 Qualitative analysis
TABLE III: Comparison with state-of-the-art methods on the test set of the LoveDA dataset. Per-class best performance is marked in bold. The FLOPs is calculated with the size of $512\times 512\times 3$. Please note that the LoveDA dataset requires an online test to evaluate the model. Therefore results for the F1 and mAcc metrics are not available here. The best value in each column is bolded.
@@ V-D1 Analysis of the model structure
TABLE V: Ablation Study on the Impact of Different ATB Strategies on the Vaihingen Dataset. The first line is base. Green and blue numbers represent increases and decreases compared to base, respectively.
@@
V-D2 Analysis of the ATB design
We also conduct experiments on the Vaihingen dataset to verify the effectiveness of the ATB design. Specifically, we use the original ATB as the baseline and remove the scaling factor $\Psi$, rotation factor $\Theta$ and offset factor $\Delta$, respectively. The experimental results are shown in Table V. When any one of these factors is removed, the performance of the model is degraded to some extent. In particular, when the scaling factor $\Psi$ is removed, the mIoU metric of the model decreases significantly. This can be explained by the fact that remote sensing images tend to have large scale variations, and a fixed local window cannot adapt to target objects of different scales. In addition, when the rotation and offset factors are removed, our model’s ability to model objects with different orientations and positions decreases, which lead to the quality of the extracted local class centers decreased. As a result, the segmentation performance of LOGCAN++ is also degraded.
V-D3 Analysis of the patch number
We conduct experiments on the Vaihingen dataset to verify the effect of different number of local patches on the results (shown in Table VI). Intuitively, the number of patches affects the size of the local window and thus the generation of local class centers. Since we split along the width and height of the feature map separately, we denote the patch number in the form of $N_{h}\times N_{w}$, e.g., $2\times 2$ means that we split the original feature map into four patches, while $1\times 1$ means that no split operation is performed, i.e., the global class center is used for context modeling in LCA module. It can be observed that as the number of patches increases, the performance of the model shows a general trend of increasing and then decreasing, and achieves an optimal value of 74.72% at $4\times 4$. Therefore, we choose the number of patches to be $4\times 4$ in the final model.
TABLE VI: Ablation Study on the Impact of Patch Numbers (i.e., $N_{h}\times N_{w}$) on the Vaihingen Dataset. The highest value is bolded.
@@ V-D3 Analysis of the patch number
$8\times 8$ 84.67 90.59 69.55 79.0 76.66 37.58 73.01
$16\times 16$ 85.07 91.46 69.82 78.97 75.99 33.46 72.46
TABLE VII: Ablation Study on the Impact of Head Numbers on the Vaihingen Dataset. The highest value is bolded.

### v3

# LOGCAN++: ADAPTIVE LOCAL-GLOBAL CLASS-AWARE NETWORK FOR SEMANTIC SEGMENTATION OF REMOTE SENSING IMAGES

@@ Abstract
Remote sensing images usually characterized by complex backgrounds, scale and orientation variations, and large intra-class variance. General semantic segmentation methods usually fail to fully investigate the above issues, and thus their performances on remote sensing image segmentation are limited. In this paper, we propose our LOGCAN++, a semantic segmentation model customized for remote sensing images, which is made up of a Global Class Awareness (GCA) module and several Local Class Awareness (LCA) modules. The GCA module captures global representations for class-level context modeling to reduce the interference of background noise. The LCA module generates local class representations as intermediate perceptual elements to indirectly associate pixels with the global class representations, targeting at dealing with the large intra-class variance problem. In particular, we introduce affine transformations in the LCA module for adaptive extraction of local class representations to effectively tolerate scale and orientation variations in remote sensing images. Extensive experiments on three benchmark datasets show that our LOGCAN++ outperforms current mainstream general and remote sensing semantic segmentation methods and achieves a better trade-off between speed and accuracy.
@@ I Introduction
Semantic segmentation of remote sensing images, aims to assign explicit categories to each pixel of a geospatial object in an image. This technique has a wide range of applications in fields such as environmental protection [1], resource investigation [2], and urban planning [3, 4]. In recent years, the rapid development of remote sensing platforms and sensors has resulted in higher spatial, spectral and temporal resolution of remote sensing images [5]. This provides the possibility of detailed observation of the Earth’s surface features. As a result, more advanced semantic segmentation techniques for remote sensing images are constantly developed to meet the needs to process rich data [6, 7].
@@
In this paper, we focus on designing a effective network structure to address the aforementioned issues. We first analyze the convolution-based neural networks. Existing convolutional neural network (CNN)-based semantic segmentation methods focus on context modeling [8, 9, 10, 11, 12], which can be divided into spatial context modeling and relational context modeling. Spatial context modeling methods, such as PSPNet [8] and DeepLabv3+ [13], use spatial pyramid pooling (SPP) or empty spatial pyramid pooling (ASPP) to integrate spatial context information. Although these methods can capture homogeneous contextual dependencies, they ignore the differences between classes, which may produce unreliable contextual information when dealing with remote sensing images with complex features and large spectral differences. Relational context modeling methods capture heterogeneous dependencies based on attention mechanism. nonlocal neural network [9] is proposed to compute the similarity of pixel pairs within an image for weighted aggregation, while DANet [10] uses spatial attention and channel attention for selective aggregation. However, such intensive attention operations are not applicable to remote sensing images, as their complex backgrounds would cause these operations to generate a large amount of background noises, leading to degraded semantic segmentation performances. Although recently proposed class-level context modeling approaches (e.g., ACFNet [14], OCRNet [15], CCENet [16] and HMANet [17]) can reduce background noise interference caused by dense attention by integrating class-level contexts (i.e., class centers) using global class representations, they are still not perform well in remote sensing images. This is caused by the fact that remote sensing images have large intra-class variance, which leads to a large semantic gap between features and global class centers, and thus hindering models’ performances.
@@
  • We introduce the affine transformation block for the first time to extract local class centers to adapt to geospatial objects with different sizes, shapes and orientations. This can cope well with the scale and orientation variations of remote sensing images.
@@
This paper is an extension of our previous conference paper [23]. This work extends the previous work in three aspects. 1) We introduce affine changes to improve the LCA module. It copes well with scale and orientation variations of remote sensing images by converting the default local window to a target quadrilateral to accommodate geospatial objects of different sizes, shapes, and orientations. 2) We extend the attention operation in the LCA module to be multi-head. so as to improve its ability to capture complex pixel-class-center relationships and to help avoid the overfitting problem that may result from a single attention mechanism. 3) We standardized the experimental setup and added sufficient experiments. We first added experiments on the larger LoveDA dataset. In addition, we provide more quantitative comparison results, qualitative analysis, and ablation experiments. We demonstrate the effectiveness of LOGCAN++ with more adequate experimental data.
@@ II-A General semantic segmentation
In summary, the above methods have achieved remarkable results on natural images. However, when applied to remote sensing images, they perform unsatisfactorily by ignoring the characteristics of remote sensing images with complex backgrounds, scale and orientation variations, and high intra-class variance. For example, recent general segmentation methods such as Segmenter [18] and DDP [48], which possesses a larger number of parameters and computational consumption, have not shown performance on remote sensing images that matches the model size (details can be found in Table I, II and III). This motivates us to propose LOGCAN++, whose network structure is designed to target these properties of remote sensing images.
@@ II-B Semantic segmentation of remote sensing images
Recently, segmentation models for generic application scenarios (e.g., LANet[59], MANet [60], PointFlow[61], SCO[62], GLOTs [63] and Farseg++[64]) also have been widely investigated. However, these methods still fail to fully consider the characteristics of remote sensing images with complex backgrounds, scale and orientation variations, and high intra-class variance, limiting the further improvement of model performance. For example, Ding et al. [59] proposes to enhance the embedding of contextual information using local attention blocks, which works to mitigate object scale variation and intra-class variance but ignores the challenges posed by complex backgrounds. Zheng et al. [65, 64] enhances the differentiation of foreground features by learning foreground-scene relations to associate foreground-related contexts. Although it enhances the features in the foreground well, it ignores the large intraclass variance in remote sensing images, leading to poorer intraclass feature compactness. Especially, these methods have insufficient consideration for the characteristics of scale and orientation variations in remote sensing images. Therefore, the performance of these methods still needs to be improved due to the deficiencies in structural design.
@@ III-C Local Class Aware Module
After the local class center $C_{l}$ is obtained, we implement the interaction between features and class centers using the multi-head cross-attention operation. Specifically, for each head $H_{i}$, the corresponding affinity matrix $R_{r}^{i}$, which represents the similarity between the pixel and the local class representations, is obtained as follows,
@@ III-D Affine Transformation Block
The ideal local class center should be the average semantic feature of objects in the local area. However, due to the Satellite sensor’s unique overhead view and large imaging range, objects in remote sensing images usually present different orientations and possess significant scale differences. Fixed local windows often do not model these objects well. Therefore, we propose an affine transformation block (ATB) to extract higher quality local class centers, where the fixed local window is converted to a target quadrilateral to accommodate geospatial objects with different sizes, shapes, and orientations.
@@
Then, we obtain the scaling factor $\Psi$ that resizes the local window, rotation factor $\Theta$ that orients the local window, and offset factor $\Delta$ that repositions the local window by feeding the obtained local contextual feature $G_{l}$ to linear mapping and LeakyReLU activation function as:
$\Theta,\Delta,\Psi=\text{LeakyReLU}(\phi_{l}(G_{l})).$ (7)
where $\phi_{l}$ denotes the linear mapping operation that outputs three representations from the obtained $G_{l}$, which are undergo the LeakyReLU activation to generate the target scaling factor $\Psi$, rotation factor $\Theta$ and offset factor $\Delta$ of the corresponding local window. Then, we give a brief description of the affine transformation block. As shown in the lower right of Fig. 3, given an initial local window, $x_{l}$, $y_{l}$ and $x_{r}$, $y_{r}$ denote the coordinates of the upper-left and lower-right corners of the initial window, respectively, while $x_{c}$, $y_{c}$ denote the coordinates of the window center. Then, its window transformation process is defined as:
$\begin{bmatrix}\hat{x}_{l/r}\\ \hat{y}_{l/r}\end{bmatrix}=\begin{bmatrix}x_{c}\\ y_{c}\end{bmatrix}+\begin{bmatrix}\Delta_{x}\\ \Delta_{y}\end{bmatrix}+\begin{bmatrix}\cos\Theta&\sin\Theta\\ -\sin\Theta&\cos\Theta\end{bmatrix}\begin{bmatrix}|x_{l/r}-x_{c}|\cdot\Psi\\ |y_{l/r}-y_{c}|\cdot\Psi\end{bmatrix},$ (8)
where $\hat{x}_{l}$, $\hat{y}_{l}$ and $\hat{x}_{r}$, $\hat{y}_{r}$ denote the coordinates of the upper-left and lower-right corners of the transformed window, respectively. $\Delta_{x}$ and $\Delta_{y}$ denote the offset in the horizontal and vertical directions, respectively. We then follow this formula to scale, rotate and offset the initial window to obtaind transformed local featurte $\hat{R}_{l}$ and corresponding pre-classified distribution $\hat{D}_{l}$. Since these factors are learned based on the pixel features, the transformed local window can be adaptively adjusted according to the size and orientation of the feature object. As a result, our ATB can help to cope with the scale and orientation changes in the remote sensing image well.
@@ IV-B Evaluation Metrics
TABLE II: Comparison with state-of-the-art methods on the test set of the ISPRS Postdam dataset. Per-class best performance is marked in bold. The best value in each column is bolded.
@@
IV-C Implementation Details
V EXPERIMENTAL RESULTS AND ANALYSIS
V-A Results on the ISPRS Vaihingen dataset
V-A1 Quantitative analysis
To validate the effectiveness of our LOGCAN++, we first carry out experiments on the Vaihingen dataset. The comparison methods include convolution-based methods such as PSPNet [8], DeepLabV3+ [13], Semantic FPN [67], DANet [10], OCRNet [15], FarSeg [65], transformer-based methods such as segmenter [18], DC-Swin [68], UNetFormer [20], EfficientViT [69] and diffusion based methods such as DDP [48]. As shown in Table I, our LOGCAN++ achieved the optimal segmentation performance. Specifically, compared to recent general segmentation methods such as DDP and RS segmentation methods such as UNetFormer, we have an increase of 2.05% and 1.23% in mIoU, respectively. In particular, for classes with large variance such as buildings, our approach achieved a mIoU value of 91.85%, which outperforms other mainstream methods, validating the effectiveness of our local global class aware strategy. In addition, for small objects such as car, we have a 1.82% mIoU improvement compared to DDP. This indicates that the introduction of affine transformations enhanced the fine-grained recognition of small objects.
In addition, LOGCAN++ possesses good efficiency compared to a series of recent popular works. Specifically, we first remove the GCA module and replace it with local class centers, resulting in a decrease of 0.1M in model parameters and 0.01G in FLOPs. Next, we eliminate the LCA module and substitute the local class centers with global class centers, which leads to a reduction of 2.06M in model parameters and 10.744G in FLOPs. These results demonstrate the lightweight of the LCA and GCA module designs. When compared to the recent state-of-the-art work DDP, LOGCAN++ requires only 78% parameters and 21% FLOPs of DDP, but exhibits superior performance (i.e., 84.22 in F1). Compared to the lightweight methods UNetFormer and EfficientViT, LOGCAN++ has similar or slightly better efficiency, yet achieves a significant performance improvement, with 1.52 and 7.78 improvements in F1 metrics, respectively. Overall, the proposed LOGCAN++ achieve a better trade-off between accuracy and efficiency.
@@ V-A2 Qualitative analysis
In addition, Fig. 5 visualizes the segmentation masks output by different models to qualitatively compare the segmentation performance of our LOGCAN++ with these competitors, where all input images come from the Vaihingen test set. Specifically, for the first input image, UNetFormer [20] and EfficientViT [69] suffer from mask boundary blurring on the building class. For the second input image, FarSeg [65] suffers from fragmentation in the building class mask, i.e., some pixels are misclassified as impervious surfaces. In contrast to these works, buildings have more complete shapes and clearer boundaries for both images in the output mask of LOGCAN++. For the third image, our LOGCAN++ is able to recognize cars accurately, while other methods such as UNetFormer and DDP [48] misidentify them as buildings. In addition, LOGCAN++ provides a complete segmentation of low-vegetation areas, whereas FarSeg appears to have some pixels that are misclassified as impervious surfaces due to these pixels having very similar spectra and shapes with impervious surfaces. These results show that our LOGCAN++ has better semantic recognition ability and visualization performance.
TABLE III: omparison with state-of-the-art methods on the test set of the LoveDA dataset. Per-class best performance is marked in bold. Please note that the LoveDA dataset requires an online test to evaluate the model. Therefore results for the F1 and mAcc metrics are not available here. The best value in each column is bolded.
@@
V-B Results on the ISPRS Potsdam dataset
V-B1 Quantitative analysis
V-B2 Qualitative analysis
@@ V-D1 Analysis of the model structure
TABLE V: Ablation Study on the Impact of Using Different Layer Features for calculating Global Class Center $C_{g}$ on the Vaihingen Dataset. The resolution increases from top to bottom. Green and blue numbers represent increases and decreases compared to base, respectively.
generation of $C_{g}$ Impervious surface Building low vegetation Tree Car Clutter mIoU
$feature_{4}$ 85.87 91.85 71.44 80.35 80.14 38.63 74.72
[…]
V-D2 Analysis of the generation of the global class center $C_{g}$
Table V compares the model performance achieved by applying four feature maps (of different resolutions) generated by four different layers of the backbone network to perform pre-classification and derive global class centers. Our findings reveal that when the model uses the feature map of the deepest layer, the model performance reaches its peak, which particularly show large performance improvements for small objects such as cars. We argue that the deepest feature map effectively eliminates much of the noise interference, allowing the model to focus on the critical information of small target objects and thereby enhancing its recognition capability.
TABLE VI: Ablation Study on the Impact of Different ATB Strategies on the Vaihingen Dataset. The first line is base. Green and blue numbers represent increases and decreases compared to base, respectively.
@@ V-D2 Analysis of the generation of the global class center $C_{g}$
[…]
TABLE VII: Ablation Study on the Impact of Patch Numbers (i.e., $N_{h}\times N_{w}$) on the Vaihingen Dataset. The highest value is bolded.
@@
[…]
V-D3 Analysis of the ATB design
We also conduct experiments on the Vaihingen dataset to verify the effectiveness of the ATB design. Specifically, we use the original ATB as the baseline and remove the scaling factor $\Psi$, rotation factor $\Theta$ and offset factor $\Delta$, respectively. The experimental results are shown in Table VI. When any one of these factors is removed, the performance of the model is degraded to some extent. In particular, when the scaling factor $\Psi$ is removed, the mIoU metric of the model decreases significantly. This can be explained by the fact that remote sensing images tend to have large scale variations, and a fixed local window cannot adapt to target objects of different scales. In addition, when the rotation and offset factors are removed, our model’s ability to model objects with different orientations and positions decreases, which lead to the quality of the extracted local class centers decreased. As a result, the segmentation performance of LOGCAN++ is also degraded. In addition, we specifically evaluate the contribution of the proposed affine transformer. Table VI compares the results achieved by our approach with and without the affine transform block (i.e., we remove all three affine transform factors). It can be observed that when the affine transform block is removed, the model performance decreases by 1.70 F1 value compared to the baseline and is lower than when only one affine transform factor is removed. The combined results show that ATB effectively improves the model segmentation performance.
V-D4 Analysis of the patch number
We conduct experiments on the Vaihingen dataset to verify the effect of different number of local patches on the results (shown in Table VII). Intuitively, the number of patches affects the size of the local window and thus the generation of local class centers. Since we split along the width and height of the feature map separately, we denote the patch number in the form of $N_{h}\times N_{w}$, e.g., $2\times 2$ means that we split the original feature map into four patches, while $1\times 1$ means that no split operation is performed, i.e., the global class center is used for context modeling in LCA module. It can be observed that as the number of patches increases, the performance of the model shows a general trend of increasing and then decreasing, and achieves an optimal value of 74.72% at $4\times 4$. Therefore, we choose the number of patches to be $4\times 4$ in the final model.
TABLE VIII: Ablation Study on the Impact of Head Numbers on the Vaihingen Dataset. The highest value is bolded.